OpenStudy (nvafer):

Show that cos(wt-b), cos(wt) and sin(wt) are functions of "t" linearly independent

OpenStudy (bonnieisflash1.0):

need help @nvafer

OpenStudy (nvafer):

Yes please @bonnieisflash1.0

OpenStudy (bonnieisflash1.0):

okay i can help

OpenStudy (bonnieisflash1.0):

what do you mean with your question

OpenStudy (nvafer):

I have to find a way to demonstrate that cos(wt-b), cos(wt) and sin(wt) are functions of "t" linearly independent i.e. that i can express one interms of other one

OpenStudy (bonnieisflash1.0):

got a website it can help

OpenStudy (bonnieisflash1.0):

that's a lot i know but don't read the whole thing

OpenStudy (bonnieisflash1.0):

i can't give answers away help you so you can understand

OpenStudy (nvafer):

It's the same my book has but i dont get it, maybe using some trigonometrical function

OpenStudy (bonnieisflash1.0):

maybe

OpenStudy (nvafer):

I know well thanks anyways

OpenStudy (bonnieisflash1.0):

your welcome

OpenStudy (bonnieisflash1.0):

still need help tho

OpenStudy (bonnieisflash1.0):

i can always help

OpenStudy (bonnieisflash1.0):

well bye @nvafer

OpenStudy (jango_in_dtown):

check the Wronskian.

OpenStudy (jango_in_dtown):

@nvafer

OpenStudy (nvafer):

@jango_IN_DTOWN im working on that thanks

OpenStudy (jango_in_dtown):

the wronskian is zero so the given quantities are linearly dependent

OpenStudy (518nad):

hi lienarly independant means u cant express in terms of the otehr

OpenStudy (holsteremission):

\[\cos(\omega t-b)=\cos\omega t\cos b+\sin\omega t\sin b\]where \(\cos b\) and \(\sin b\) are presumably constant. So immediately you can see that the first function is a linear combination of the other two.